OpenStudy (anonymous):

What is the simplified form of 3/2x-5+21/8x^2-14x-15? Please help.

OpenStudy (anonymous):

is it \[\frac{3}{2x+5}+\frac{21}{8x^2-14x-15}\] like that ?

OpenStudy (anonymous):

Yea but its 2x-5

OpenStudy (anonymous):

oh good luckily the second denominator factors as \[(2x-5)(4x+3)\] so you can use that as a common denominator

OpenStudy (anonymous):

you get \[\frac{3(4x+3)+21}{(2x-5)(4x+3)}\] then some algebra to clean up the numerator you good from there ?

OpenStudy (anonymous):

Yes thanks!

OpenStudy (anonymous):

yw